# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v2

@@ II.1 Expansion in energy eigenbasis
In order to phrase the problem quantum-mechanically, we promote the variable $Q$ to be an operator $\hat{Q}$ in the Heisenberg picture, undergoing unitary time evolution as $\hat{Q}(t)=U^{\dagger}\hat{Q}U$, where $U=\text{exp}(-itH)$ with a standard QHO Hamiltonian $H$. We can write $\hat{Q}=(2\hat{\theta}-1)$, where $\hat{\theta}$ is the Heaviside operator. For convenience, we will work exclusively with units chosen such that $\hbar=1$; further, it will be convenient to normalise the angular frequency of oscillations to unity, $\omega=1$. The Tsirelson inequality is then $1\leq\langle A\rangle\leq 1$, where now
$A\equiv\hat{Q}+U^{\dagger}\hat{Q}U+(U^{\dagger})^{2}\hat{Q}U^{2}\ ,$ (2.1)
@@
with $U=\text{exp}(-\frac{2\pi}{3}iH)$. We will refer to $A$ hereafter as the “Tsirelson operator”. Quantum systems can violate this inequality—up to a maximal value of $\langle A\rangle\approx 1.26$, so Tsirelson found [1].
@@
We can compute the Tsirelson quantity for an arbitrary state of the QHO Hilbert space formed as a superposition of energy eigenstates from the ground state upwards up to a finite cut-off:
@@
The single-time average $\langle\hat{Q}\rangle=\langle\Psi|2\hat{\theta}-1|\Psi\rangle$ can be written as
$\langle\Psi|2\hat{\theta}-1|\Psi\rangle=\sum_{n,k=0}^{N}c_{n}^{*}Q_{nk}c_{k}\ ,$ (2.3)
where we have defined the matrix elements $Q_{nk}=\langle n|2\hat{\theta}-1|k\rangle$. A tractable expression for $Q_{nk}$ is derived in Appendix C and is given by
@@
In the Schrödinger picture,
$|\Psi,t\rangle=\sum_{n=0}^{N}c_{n}e^{-int}|n\rangle\ ,$ (2.5)
since $|n\rangle$ is by definition an eigenstate of the QHO Hamiltonian with eigenvalue $(n+\frac{1}{2})$. We can drop the vacuum contribution of $\frac{1}{2}$ since this will have no effect upon the expectation value:
$\langle\hat{Q}(t)\rangle=\langle\Psi,t|\hat{Q}|\Psi,t\rangle=\sum_{n,k=0}^{N}c_{k}^{*}c_{n}e^{-i(n-k)\omega t}Q_{nk}\ .$ (2.6)
@@
$\langle A\rangle=\sum_{n,k=0}^{N}c_{k}^{*}c_{n}(1+e^{-i\frac{2\pi}{3}(n-k)}+e^{-i\frac{4\pi}{3}(n-k)})Q_{nk}\ ,$ (2.7)
recalling that, by choosing units of $\omega=1$, one full period corresponds to $T=2\pi$.
Violations of the Tsirelson inequality are manifestly associated with negativity in the Wigner function. A corollary is that a single coherent state, which has a completely positive Wigner function, will not violate the Tsirelson inequality; neither will a single energy eigenstate, for it is trivially of strictly odd or even parity. A neat proof, in terms of probability currents, that definite-parity states do not violate the Tsirelson inequality is provided in Section IV.1. Perhaps more surprising is that a superposition of at least three energy eigenstates is required to produce a violation. We provide a full proof of this fact in Appendix E.
II.2 Maximally violating eigenstates
The Tsirelson operator is Hermitian ($\hat{A}\equiv\hat{Q}_{1}+\hat{Q}_{2}+\hat{Q}_{3}=\hat{A}^{{\dagger}}$), and so its eigenvalues $\lambda_{i}\in\mathbb{R}$ (i=0,1,,Ni=0,1,...,N) for a truncated $(N+1)$-dimensional subspace of the Hilbert space are real. The greatest eigenvalue $\lambda_{\text{max}}=\text{max}(\{\lambda_{i}\})$ corresponds to the largest possible Tsirelson violation, $\langle A\rangle>1$, for a superposition $|\psi\rangle=\sum_{n=0}^{N}c_{n}|n\rangle$ of QHO energy eigenstates up to $|N\rangle$. Likewise, the lowest eigenvalue corresponds to the most severe violation of the classical lower bound, $\langle A\rangle=-1$, in the subspace. The matrix elements of $A$ in the energy eigenbasis can be written in terms of the familiar $Q_{nk}$:
${A}_{nk}\equiv\langle n|{A}|k\rangle=Q_{nk}(1+e^{i\tau(n-k)}+e^{2i\tau(n-k)})\ ,$ (2.8)
@@ II.2 Maximally violating eigenstates
By explicitly diagonalising the Tsirelson operator $A$ in an increasingly large subspace, which is detailed in Appendix D, we find that the maximally-violating state in the $N=6$ subspace is given by:
@@
with a lower-bound Tsirelson violation $\langle A\rangle=-1.1195$. This agrees exactly with the work of Ref. [6], for which we label Eq. 2.9 the Zaw–Scarani state. We shall use this as a simple candidate state for later analysis.
@@
$|\Psi\rangle=a_{0}|0\rangle+c_{a}|3\rangle+a_{6}|6\rangle\ ,\ \text{where}\ a_{0},a_{3},a_{6}\in\mathbb{R}\ .$ (2.11)
Clearly, there exists a sizeable region of states over which a Tsirelson violation can be achieved to varying extent. This is good news for experimental realisations of the protocol, because it means even quantum states that are prepared imperfectly can, in principle, produce demonstrable Tsirelson violation. Another important observation is that every upper-bound-violating state is paired with a commensurate lower-bound-violating state ($A\rightarrow-A$), which is obtained by swapping the sign of the odd-parity coefficients (e.g. $c_{3}\rightarrow-c_{3}$).
Figure 4: Representation of all states in the subspace described by Eq. 2.11, where the axes represent the (real) coefficients. (a) A colour map is used to represent the value of the Tsirelson quantity when each of these states is taken to be the initial state. The Zaw–Scarani state, labelled $|\Psi_{Z}\rangle$, is the maximally violating state and lies in the centre of the upper-bound-violating region. (b) Only the violating (non-classical) regions are highlighted, showing that each upper violation is paired with a lower violation.
II.3 Structure of the Tsirelson operator
This section establishes a link between the Tsirelson operator and the group $C_{3}$. We then apply this connection to find an alternative explanation for a result previously noted in Ref. [4].
The group $C_{3}$ has three elements which we denote $\{e,r,r^{2}\}$. It is an Abelian group whose generator is $r$ [63]. From the Cayley table of the group, outlined in Table 1, it is straightforward to determine that $\{\mathds{1},U,U^{2}\}$ forms a faithful representation of $C_{3}$ under multiplication with $U$ acting as the generator.
@@ II.3 Structure of the Tsirelson operator
$\operatorname{Twirl}(\hat{Q})=\frac{1}{3}A=\frac{1}{3}\left[\mathds{1}Q\mathds{1}+U^{\dagger}QU+(U^{2})^{\dagger}Q(U^{2})\right]\ .$ (2.12)
The $\operatorname{Twirl}(\cdot)$ operation is the averaging of a matrix over the actions of a group, in this case $C_{3}$ [64]. It is a useful operation because it extracts the component of the matrix that transforms under the identity (often dubbed the “trivial”) representation of the group. Note that $U^{\dagger}\operatorname{Twirl}(Q)U=(1)\operatorname{Twirl}(Q)\implies[\operatorname{Twirl}(Q),U]=0$.
@@
[…]
Figure 5: The Tsirelson operator in the energy eigenbasis with any finite energy cut-off can be interpreted as the adjacency matrix of a simple weighted graph. This is illustrated with a cut-off at $n=29$, where the nodes are energy eigenstates and the edges represent non-zero matrix elements of the Tsirelson operator. The colour represents the value of the matrix element.
III Assessing the Dynamical Assumption
III.1 General Framework
We now come to the main issue, which is assessing the degree to which the dynamical assumptions behind the Tsirelson inequality hold. As discussed, like many other classical phenomena, uniform precession and its closely-related assumptions have no unique definition in quantum theory. We will therefore explore a number of inequivalent implementations which entail different measurement techniques to check it. Specifically, we will implement the following procedure in a number of different ways:
@@ III.1 General Framework
    where the scaled Tsirelson quantity $\frac{1}{2}(1+\langle A\rangle)$ is bounded between 0 and $1$ classically and hence describes a probability.
@@
III.2 Method 1: Macrorealistic Description of Sign Changes
@@
Step 2: We find, using the probability moment expansions in Eqs. 1.7 and 1.10, that the Tsirelson quantity can be written as
@@
From Eq. 3.6 and Eq. 3.7, we see that the Tsirelson violation, $\frac{1}{2}(1+\langle A\rangle)$, will itself respect the Lüders bound if uniform precession is satisfied exactly ($L_{1}=0$), but in general can exceed this bound, as is the case with the maximally-violating state in the full infinite-dimensional QHO Hilbert space, which Tsirelson found to be $\frac{1}{2}(1+\langle A\rangle)=-0.13<-0.125$ [1].
@@
The interference term in Eq. 3.1 is purely quantum-mechanical in nature; however, an important problem to consider is whether any further non-classicality is indicated by $L_{1}$, which should be zero classically. Non-classicality in temporal correlations can stem from two separate sources: non-commutativity of successive measurements and negativity of the Wigner function [65]. Whilst violations of the Tsirelson inequality require negativity of the Wigner function, the non-zero positive $L_{1}$ is an independent diffraction-in-time effect, as characterised in Ref. [66] but first studied in Refs. [67, 68]. Therefore, even a state with positive Wigner function, which will not violate the Tsirelson quantity, can render this LG3 non-zero.
More speculatively, we also note that the size of the uniform precession violation here is roughly $4\%$, which is peculiarly similar to the estimated bound on the so-called quantum back-flow effect, the phenomenon in which a particle prepared with strictly positive momentum can still exhibit negative probability current [69, 70, 17, 71].
We also note, at this point, that simple superpositions of coherent states—such as the “three-headed cat state”—will violate the Tsirelson quantity [5]. We have also verified in our calculations that the state
@@
with the appropriate normalisation factor $N$, yields a Tsirelson violation of $\langle A\rangle=-1.0795<-1$. Thus, the method of assessing UP with an LG quantity can also be used upon such states, which may be of greater interest experimentally, for Gaussian state superpositions can be easier to create in a laboratory setting [72].
III.3 Method 2: Three-time Sequential Measurements
Our second method uses a generalised dynamical assumption involving $\Delta p=p(+,+,+)-p(-,-,-)$, where $p$ is a probability to be specified. This is of interest because, classically, the Tsirelson quantity $A$ can take values $\pm 3$ and $\pm 1$, but setting $\Delta p=0$ prevents the values $\pm 3$ from contributing to the computation of the average $\langle A\rangle$, and hence trivially $1\leq\langle A\rangle\leq 1$.
The probabilities could be specified to involve three-time projective measurements, $\Delta p_{123}$, corresponding to Eq. 1.11, or “weak” measurements implemented at times $t_{1}$ or $t_{2}$, denoted by $\Delta p_{(1)23}$ and $\Delta p_{1(2)3}$, respectively. Finally, we could also consider the quasi-probability $\Delta q$, corresponding to Eq. 1.7. Appendix H details the computational approach to calculating all such measurement probabilities. An important point is that all of these four distributions are identical to a macrorealist but differ quantum mechanically by various interference terms.
@@ III.3 Method 2: Three-time Sequential Measurements
When weak measurements are employed, UP is satisfied only marginally in the case of $\Delta p_{1(2)3}$; and when the quasi-probability distribution $\Delta q$ is instead used, which produces no interference terms, the UP violation is exactly commensurate with the Tsirelson violation. This is to be expected, since $\Delta q=0$ is a sufficient condition to derive the Tsirelson inequality, which property follows directly from the moment expansion form of Eq. 1.7. The quasi-probabilities, though they can be negative, respect the sum rules for probabilities; and when a measurement is rendered weak, the probability sum rule holds valid over that particular measurement time, as in $p_{1(2)3}(+,+,+)+p_{1(2)3}(+,-,+)=p_{13}(+,+)$.
  Measure of UP
Interference term, II
$2\Delta p$
UP satisfied?
($2|\Delta p|<\frac{1}{2}|1+\langle A\rangle|=0.0598$ ?)
@@
$\Delta p_{123}$
$\frac{1}{2}(\langle\hat{Q}_{2}\rangle-\langle\hat{Q}_{2}^{(1)}\rangle)+\frac{1}{2}(\langle\hat{Q}_{3}\rangle-\langle\hat{Q}_{3}^{(12)}\rangle)$
0.0191-0.0191
$\Delta p_{(1)23}$ $\frac{1}{2}(\langle\hat{Q}_{3}\rangle-\langle\hat{Q}_{3}^{(2)}\rangle)$ 0.0785-0.0785
$\Delta p_{1(2)3}$
$\frac{1}{2}(\langle\hat{Q}_{2}\rangle-\langle\hat{Q}_{2}^{(1)}\rangle)+\frac{1}{2}(\langle\hat{Q}_{3}\rangle-\langle\hat{Q}_{3}^{(1)}\rangle)$
0.0573-0.0573
Table 2: The various means by which the dynamical assumption of uniform precession (UP) can be assessed through three-time probability distributions in Method 2 of Section III. A subscript in brackets indicates that the measurement at that time was weak. A tick indicates that the UP violation is too small to produce the overall violation, which must therefore be produced by the quantum-mechanical interference terms.
@@
IV Heuristic Notions of Uniform Precession

### v3

@@ III.1 Expansion in energy eigenbasis
In order to discuss the Tsirelson quantity quantum-mechanically, we promote the variable $Q$ to be an operator in the Heisenberg picture, undergoing unitary time evolution as $\hat{Q}(t)=U^{\dagger}\hat{Q}U$, where $U=\text{exp}(-itH)$ with a standard QHO Hamiltonian $H$. We can write $\hat{Q}=(2{\theta}(\hat{x})-1)$, where ${\theta}(\hat{x})=\int_{0}^{\infty}dx|x\rangle\langle x|$ is the Heaviside operator (and $|x\rangle$ are position eigenstates). For convenience, we will work exclusively with units chosen such that $\hbar=1$; further, it will be convenient to normalise the angular frequency of oscillations to unity, $\omega=1$. The Tsirelson inequality is then $1\leq\langle A\rangle\leq 1$, where now
$\hat{A}\equiv\hat{Q}+U^{\dagger}\hat{Q}U+(U^{\dagger})^{2}\hat{Q}U^{2}\ ,$ (3.1)
@@
with $U=\text{exp}(-\frac{2\pi}{3}iH)$, recalling that, by choosing units of $\omega=1$, one full period corresponds to $T=2\pi$. We will refer to $\hat{A}$ hereafter as the “Tsirelson operator”. Quantum systems can violate this inequality—up to a maximal value of $\langle A\rangle\approx 1.26$, so Tsirelson found [tsirelson_how_2006].
@@
We can compute the Tsirelson quantity for an arbitrary state of the QHO Hilbert space formed as a superposition of energy eigenstates, $|n\rangle$, from the ground state upwards up to a finite cut-off:
@@
The single-time average $\langle Q\rangle=\langle\Psi|2\theta(\hat{x})-1|\Psi\rangle$ can be written as
$\langle\Psi|2\theta(\hat{x})-1|\Psi\rangle=\sum_{n,k=0}^{N}c_{n}^{*}Q_{nk}c_{k}\ ,$ (3.3)
where we have defined the matrix elements $Q_{nk}=\langle n|2\theta(\hat{x})-1|k\rangle$. A tractable expression for $Q_{nk}$ is derived in Appendix C and is given by
@@
So, for a state in the form of Eq. 3.2,
$\langle Q(t)\rangle=\langle\Psi|\hat{Q}|\Psi\rangle=\sum_{n,k=0}^{N}c_{k}^{*}c_{n}e^{-i(n-k)t}Q_{nk}\ .$ (3.5)
@@
$\langle A\rangle=\sum_{n,k=0}^{N}c_{k}^{*}c_{n}(1+e^{-i\frac{2\pi}{3}(n-k)}+e^{-i\frac{4\pi}{3}(n-k)})Q_{nk}\ .$ (3.6)
Violations of the Tsirelson inequality are manifestly associated with negativity in the Wigner function. A corollary is that a single coherent state, which has a completely positive Wigner function, will not violate the Tsirelson inequality; neither will a single energy eigenstate, for it is trivially of strictly odd or even parity. A proof, in terms of probability currents, that definite-parity states do not violate the Tsirelson inequality is provided in Section V.1. Perhaps more surprising is that a superposition of at least three energy eigenstates is required to produce a violation. We provide a full proof of this fact in Appendix E.
III.2 The simplest violating state
The Tsirelson operator is Hermitian ($\hat{A}\equiv\hat{Q}_{1}+\hat{Q}_{2}+\hat{Q}_{3}=\hat{A}^{{\dagger}}$), and so its eigenvalues for a truncated $(N+1)$-dimensional subspace of the Hilbert space are real. The greatest eigenvalue corresponds to the largest possible Tsirelson violation, $\langle A\rangle>1$, for a superposition $|\psi\rangle=\sum_{n=0}^{N}c_{n}|n\rangle$ of QHO energy eigenstates up to $|N\rangle$. Likewise, the lowest eigenvalue corresponds to the most severe violation of the classical lower bound, $\langle A\rangle=-1$, in the subspace. The matrix elements of $\hat{A}$ in the energy eigenbasis can be written in terms of the familiar $Q_{nk}$:
${A}_{nk}\equiv\langle n|{A}|k\rangle=Q_{nk}(1+e^{i\frac{2\pi}{3}(n-k)}+e^{i\frac{4\pi}{3}(n-k)})\ ,$ (3.7)
@@ III.2 The simplest violating state
By explicitly diagonalising the Tsirelson operator $\hat{A}$ in an increasingly large subspace, which is detailed in Appendix D, we find that the greatest Tsirelson violation in the $N=6$ subspace is given by:
@@
with a lower-bound Tsirelson violation $\langle A\rangle=-1.1195$. This agrees exactly with the work of Ref. [PhysRevA.108.022211], for which we label Eq. 3.8 the Zaw–Scarani state. We shall use this as a simple candidate state for later analysis because the $N=6$ cut-off is the first point at which violations to Tsirelson’s inequality become possible (a direct consequence of Appendix E and Section III.3).
@@
$|\Psi\rangle=a_{0}|0\rangle+a_{3}|3\rangle+a_{6}|6\rangle\ ,\ \text{where}\ a_{0},a_{3},a_{6}\in\mathbb{R}\ .$ (3.10)
Clearly, there exists a sizeable region of states over which a Tsirelson violation can be achieved to varying extent. This is good news for experimental realisations of the protocol, because it means even quantum states that are prepared imperfectly can, in principle, produce demonstrable Tsirelson violation. Another important observation is that every upper-bound-violating state is paired with a commensurate lower-bound-violating state ($\langle A\rangle\rightarrow-\langle A\rangle$), which is obtained by swapping the sign of the odd-parity coefficients (e.g. $c_{3}\rightarrow-c_{3}$).
Figure 4: Representation of all states in the subspace described by Eq. 3.10, where the axes represent the (real) coefficients. (a) A colour map is used to represent the value of the Tsirelson quantity when each of these states is taken to be the initial state. The Zaw–Scarani state, labelled $|\Psi_{Z}\rangle$, has the greatest violation in this subspace and lies in the centre of the upper-bound-violating region. (b) Only the violating (non-classical) regions are highlighted, showing that each upper violation is paired with a lower violation.
III.3 Structure of the Tsirelson operator
This section establishes a link between the Tsirelson operator and the group $C_{3}$. We then apply this connection to find an alternative explanation for a result previously noted in Ref. [precessions].
The group $C_{3}$ has three elements which we denote $\{e,r,r^{2}\}$. It is an Abelian group whose generator is $r$ [groupC3]. From the Cayley table of the group, outlined in Table 1, it is straightforward to determine that $\{\mathds{1},U,U^{2}\}$ forms a faithful representation of $C_{3}$ under multiplication with $U$ acting as the generator.
@@ III.3 Structure of the Tsirelson operator
$\operatorname{Twirl}(\hat{Q})=\frac{1}{3}\hat{A}=\frac{1}{3}\left[\mathds{1}\hat{Q}\mathds{1}+U^{\dagger}\hat{Q}U+(U^{2})^{\dagger}\hat{Q}(U^{2})\right]\ .$ (3.11)
The $\operatorname{Twirl}(\cdot)$ operation is the averaging of a matrix over the actions of a group, in this case $C_{3}$ [twirling]. It is a useful operation because it extracts the component of the matrix that transforms under the identity representation of the group. Note that $U^{\dagger}\operatorname{Twirl}(\hat{Q})U=(1)\operatorname{Twirl}(\hat{Q})\implies[\operatorname{Twirl}(\hat{Q}),U]=0$.
@@
Therefore, the eigenvectors of $\operatorname{Twirl}({Q})$, and hence $A$, must be made up exclusively of basis vectors in a single degenerate subspace. This property is illustrated in Figure 5 by representing the Tsirelson operator as a graph, where each subgraph corresponds to one of the subspaces. The bipartite nature of the subgraphs visualises the underlying symmetric spectrum of the operator.
Figure 5: The Tsirelson operator in the energy eigenbasis with any finite energy cut-off can be interpreted as the adjacency matrix of a simple weighted graph. This graph is illustrated with a cut-off at $n=29$, where the nodes are energy eigenstates and the edges represent non-zero matrix elements of the Tsirelson operator. The colour represents the value of the matrix element.
IV Assessing the Dynamical Assumption
IV.1 General framework
We now come to the main issue, which is assessing the degree to which the dynamical assumptions behind the Tsirelson inequality hold. As discussed, like many other classical phenomena, uniform precession and its closely-related assumptions have no unique definition in quantum theory. We will therefore explore a number of inequivalent implementations which entail different measurement techniques to check it. Our first method involves demonstrating an explicit connection between $\langle A\rangle$ and the probability of zero sign changes, L1 (Eqs. 2.9 and 2.13), which should be zero classically. Our second method stems from the interesting observation that a weaker condition of setting merely the difference q(+,+,+)q(,,)q(+,+,+)-q(-,-,-) to be zero also leads to the Tsirelson inequality, as we shall show, and thus this may also be considered a measure of uniform precession.
Specifically, we will implement the following procedure in a number of different ways:
@@ IV.1 General framework
    The specific form on the left-hand side is used because $\frac{1}{2}(1\pm\langle A\rangle)$ are the probabilities for $A=\pm 1$ when it can take only values $\pm 1$, as is the case classically (since uniform precession holds). This gives a convenient scale against which to compare violations of the Tsirelson inequality and of uniform precession.
@@
IV.2 Method 1: Macrorealistic description of sign changes
@@
Step 2: We find, using the probability moment expansions in Eqs. 2.3 and 2.6, that the Tsirelson quantity can be written as
@@
From Eq. 4.6 and Eq. 4.7, we see that the Tsirelson violation, $\frac{1}{2}(1+\langle A\rangle)$, will itself respect the Lüders bound if uniform precession is satisfied exactly ($L_{1}=0$), but in general can exceed this bound, as is the case with the maximally-violating state in the full infinite-dimensional QHO Hilbert space, which Tsirelson found to be $\frac{1}{2}(1+\langle A\rangle)=-0.13<-0.125$ [tsirelson_how_2006].
@@
We also note, at this point, that simple superpositions of coherent states—such as the “three-headed cat state”—will violate the Tsirelson quantity [PhysRevLett.130.160201]. For example,
@@
with the appropriate normalisation factor $N$, yields a Tsirelson violation of $\langle A\rangle=-1.0795<-1$. Thus, the method of assessing UP with an LG quantity can also be used upon such states, which may be of greater interest experimentally, for Gaussian state superpositions can be easier to create in a laboratory setting [Ourjoumtsev2007].
IV.3 Method 2: Three-time sequential measurements
Our second method uses a dynamical assumption involving $\Delta p=p(+,+,+)-p(-,-,-)$, where $p$ is a probability to be specified, which weakens the original assumption for deriving Tsirelson’s inequality [tsirelson_how_2006, plavala_tsirelson_2024, precessions]. This is of interest because, classically, the Tsirelson quantity $A$ can take values $\pm 3$ and $\pm 1$, but setting $\Delta p=0$ prevents the values $\pm 3$ from contributing to the computation of the average $\langle A\rangle$, and hence trivially $1\leq\langle A\rangle\leq 1$.
The probabilities could be specified to involve three-time projective measurements, $\Delta p_{123}$, corresponding to Eq. 2.7, or we could consider the quasi-probability $\Delta q$, corresponding to Eq. 2.3. We could also use the two quasi-probability distributions mentioned in the discussion surrounding Eq. 2.8, which lie between $\Delta p_{123}$ and $\Delta q$. We shall denote these distributions $\Delta q_{A}$, $\Delta q_{B}$, which we define fully in Appendix H). Appendix H also details the computational approach to calculating all such measurement probabilities. An important point is that all of these four distributions are identical to a macrorealist but differ quantum mechanically by various interference terms.
@@ IV.3 Method 2: Three-time sequential measurements
In the case of the distribution $\Delta q_{B}$, UP is satisfied only marginally; and when the quasi-probability distribution $\Delta q$ is instead used, which produces no interference terms, the UP violation is exactly commensurate with the Tsirelson violation. This is to be expected, since $\Delta q=0$ is a sufficient condition to derive the Tsirelson inequality, which property follows directly from the moment expansion form of Eq. 2.3.
  Measure of UP Interference term, II $2\Delta p$ UP satisfied sufficiently? ($2|\Delta p|<\frac{1}{2}|1+\langle A\rangle|=0.0598$ ?)
@@
$\Delta p_{123}$ $\frac{1}{2}(\langle{Q}_{2}\rangle-\langle{Q}_{2}^{(1)}\rangle)+\frac{1}{2}(\langle{Q}_{3}\rangle-\langle{Q}_{3}^{(12)}\rangle)$ 0.0191-0.0191
$\Delta q_{A}$ $\frac{1}{2}(\langle{Q}_{3}\rangle-\langle{Q}_{3}^{(2)}\rangle)$ 0.0785-0.0785
$\Delta q_{B}$ $\frac{1}{2}(\langle{Q}_{2}\rangle-\langle{Q}_{2}^{(1)}\rangle)+\frac{1}{2}(\langle{Q}_{3}\rangle-\langle{Q}_{3}^{(1)}\rangle)$ 0.0573-0.0573
Table 2: The various means by which the dynamical assumption of uniform precession (UP) can be assessed through three-time probability distributions in Method 2 of Section IV. A tick indicates that the UP violation is too small to produce the overall Tsirelson violation, which must therefore be produced, in part, by the quantum-mechanical interference terms.
@@
We also note that perfect sign-operator measurements are an idealisation for any real experimental procedure and comment on the effect of smoothed projectors in Appendix K.
V Heuristic Notions of Uniform Precession